How much are you willing to pay for one share of APOL stock if the company just paid an annual dividend of $2.25, the dividends increase by 2 percent annually, and you require a return of 15 percent?
Answer #1

Ans $ 17.65

P0 = Price of Share
D1 = Current Dividend
Ke = Cost of Equity
g = growth rate
P0 = D1 / (Ke - g)
P0 = 2.2950 / (15%- 2%)
P0 = 17.65
D1 = D0* (1 + g)
D1 = 2.25* (1 + 2.00%)
D1 = 2.2950
Answer #2

Current price=D1/(Required return-Growth rate)

=(2.25*1.02)/(0.15-0.02)

which is equal to

=$17.65(Approx).